# event-ECC: Asynchronous Tracking of Events with Continuous Optimization

### v1

@@ 3.1 Preliminaries
with $\mathbf{x}_{F}(T)$ being the position of feature $F$ in $\mathcal{X}_{\mathcal{S}_{I}}$ at that time instance, and $\theta_{F}(T)$ the orientation of its neighborhood $\mathcal{N}\Big{(}\mathbf{s}(T)\Big{)}$ of size $(2N+1)\times(2N+1)$, defined by:
@@
If we now assume a Euclidean geometric transformation parameterized by the elements of the state vector defined in Eq. (1),
that is, for each pair of corresponding points $\mathbf{x}(T)~{}\in~{}\mathcal{N}\big{(}\mathbf{s}\big{)}$ and $\mathbf{x}^{\prime}(T)~{}\in~{}\mathcal{X}_{\mathcal{S}_{\mathcal{T}_{F}}}$, then $\mathbf{x}(T)=R\big{(}\theta_{F}(T)\big{)}\mathbf{x}^{\prime}(T)+\mathbf{x}_{F}(T)$, or equivalently:
@@
where the vectors $\mathbf{v}_{i},~{}~{}i=1,2$ constitute the natural basis of $\mathbb{R}^{2}$ and $\mathbf{n}=\bigl{\lfloor}\mathbf{x}\bigr{\rfloor}$ denotes the the element-wise floor of $\mathbf{x}$. Then, similar to [4], we define a density update rule using linear interpolation as follows:
@@ 4.1 Model and Template Windows
Note that those elements of the warped template vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k+1}\big{)}$ which correspond to model’s pixel with zero intensity, attain their previous values $\mathcal{T}_{F}\big{(}\mathbf{n}_{m}^{\prime}~{}T_{k}\big{)}$.
@@ 4.2 The Proposed Optimization Criterion
Solving the optimization problem is clearly not a simple task because of the nonlinearity involved in the correspondence part. This, of course, suggests that its minimization requires nonlinear optimization techniques by adopting a gradient-based approach.
To this end, let us consider the following additive updating rule for each one of the elements of the state vector:
@@
where $\boldsymbol{x}_{m}^{\prime}\big{(}\mathbf{s}_{k+1}\big{)}~{}\in~{}\mathcal{X}_{\mathcal{S}_{\mathcal{T}_{F}}}$ and the $2\times 3$ matrix $W_{m}\big{(}\mathbf{s}_{k}\big{)}$ depends only on the state vector $\mathbf{s}_{k}$.
Note that $\boldsymbol{x}_{m}^{\prime}\big{(}\mathbf{s}_{k+1}\big{)}$ is composed by two terms. The first one depends on the state vector $\mathbf{s}_{k}$ which is known, while the second one depends on the unknown perturbation vector $\Delta\boldsymbol{s}_{k+1}$. Therefore, using first order Taylor approximation around $\mathbf{s}_{k}$ for every element of the warped vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k+1}\big{)}$, we obtain:
@@
where $J\big{(}\mathbf{s}_{k}\big{)}$ denotes the $(2N+1)^{2}\times 3$ Jacobian matrix of the warped intensity vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k}\big{)}$, evaluated at the nominal parameter values $\mathbf{s}_{k}$, and $\Delta\mathbf{s}_{k+1}$ denotes the perturbations of the parameters of the state vector. By substituting this approximation into Eq. ( 12) we get the following simpler objective function:
@@ 4.3 Lightweight event-ECC
To this end, consider the set $\mathcal{N}^{\prime}_{k+1}$ with the coordinates of the four pixels which intensities were updated after the calculation of the warp, defined in Eq. (17). That is, this set contains the recently modified template’s pixels whose densities will be used in the next step of the tracking algorithm. Accordingly, we define the sets $\mathcal{N}^{\prime}_{{k+1}_{x}}$ $\mathcal{N}^{\prime}_{{k+1}_{y}}$ of the template’s gradient whose values have changed. Depending on the position of the pixels of $\mathcal{N}^{\prime}_{k+1}$ in the support area $\mathcal{S}_{T}$ of the template, the cardinality of the union $\mathcal{S}=\mathcal{N}^{\prime}_{{k+1}_{x}}\bigcup\mathcal{N}^{\prime}_{{k+1}_{y}}$ is bounded by $12$, that is $|\mathcal{S}|\leq~{}12$. Having defined the set $\mathcal{S}$ we can define its complement w.r.t. the set $\mathcal{S}_{\mathcal{T}_{F}}$, that is,
$\mathcal{S}^{c}=\mathcal{S}_{\mathcal{T}_{F}}\backslash\mathcal{S}$ (18)
@@
$J\big{(}\mathbf{s}_{k+1}\big{)}$ $=$ $\left[J_{\mathcal{S}^{c}}^{t}\big{(}\mathbf{s}_{k}\big{)}~{}J_{\mathcal{S}}^{t}\big{(}\mathbf{s}_{k+1}\big{)}\right]^{t}$
@@
$\mathbf{p}_{\mathbf{t}_{F}}(\mathbf{s}_{k+1})$ $=$ $\mathbf{p}_{\mathbf{t}_{F}}(\mathbf{s}_{k})+\Delta\mathbf{p}_{k+1}-\Delta\mathbf{p}_{k}~{}.$ (19)
@@
         Form the model’s vector $\mathbf{m}_{F}(T_{k+1})$ using the
         updated set $\mathcal{E}_{\mathcal{M}}(T_{k+1})$
         Using Eq. (17), Compute the optimal warp
         $\Delta\mathbf{s}^{\star}_{k+1}$
@@
         Use $\boldsymbol{s}_{k+1}$ in Eq. (10) to find out where the central
         event $\mathbf{x}_{c}^{\prime}(T_{k+1})$ of set $\mathcal{E}_{\mathcal{M}}(k+1)$ is mapped.
         Use the interpolation scheme (Eq. (6)) to Update the
         densities of template’s pixels $\in\mathcal{N}^{\prime}_{k+1}$
         Form the vector $\mathbf{t}_{F}(T_{k+1})$ and Compute the
         quantities of Eq. (19)
         Update the states matrix $S_{F}=[S_{F}~{}\boldsymbol{s}_{k+1}]$

### v2

@@ 3.1 Preliminaries
with $\mathbf{x}_{F}(T)$ being the 2D position of feature $F$ in $\mathcal{X}_{\mathcal{S}_{I}}$ at that time instance, and $\theta_{F}(T)$ the orientation of its neighborhood $\mathcal{N}\Big{(}\mathbf{s}(T)\Big{)}$ of size $(2N+1)\times(2N+1)$, defined by:
@@
If we now assume a Euclidean geometric transformation parameterized by the elements of the state vector defined in Eq. (1), that is, for each pair of corresponding points $\mathbf{x}(T)~{}\in~{}\mathcal{N}\big{(}\mathbf{s}\big{)}$ and $\mathbf{x}^{\prime}(T)~{}\in~{}\mathcal{X}_{\mathcal{S}_{\mathcal{T}_{F}}}$, then $\mathbf{x}(T)=R\big{(}\theta_{F}(T)\big{)}\mathbf{x}^{\prime}(T)+\mathbf{x}_{F}(T)$, or equivalently:
@@
where the vectors $\mathbf{v}_{i},~{}~{}i=1,2$ constitute the natural basis of $\mathbb{R}^{2}$ and $\mathbf{n}=\bigl{\lfloor}\mathbf{x}\bigr{\rfloor}$ denotes the element-wise floor of $\mathbf{x}$. Then, similar to [4], we define a density update rule using linear interpolation as follows:
@@ 4.1 Model and Template Windows
Note that those elements of the warped template vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k+1}\big{)}$ which correspond to model’s pixel with zero density, attain their previous values $\mathcal{T}_{F}\big{(}\mathbf{n}_{m}^{\prime}~{}T_{k}\big{)}$.
@@ 4.2 The Proposed Optimization Criterion
Solving the optimization problem is clearly not a simple task because of the nonlinearity involved in the correspondence part. This, of course, suggests that its minimization requires nonlinear optimization techniques by adopting a gradient-based approach. To this end, let us consider the following additive updating rule for each one of the elements of the state vector:
@@
where $\boldsymbol{x}_{m}^{\prime}\big{(}\mathbf{s}_{k+1}\big{)}~{}\in~{}\mathcal{X}_{\mathcal{S}_{\mathcal{T}_{F}}}$ and the $2\times 3$ matrix $W_{m}\big{(}\mathbf{s}_{k}\big{)}$ depends only on the state vector $\mathbf{s}_{k}$. Note that $\boldsymbol{x}_{m}^{\prime}\big{(}\mathbf{s}_{k+1}\big{)}$ is composed by two terms. The first one depends on the state vector $\mathbf{s}_{k}$ which is known, while the second one depends on the unknown perturbation vector $\Delta\boldsymbol{s}_{k+1}$. Therefore, using first order Taylor approximation around $\mathbf{s}_{k}$ for every element of the warped vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k+1}\big{)}$, we obtain:
@@
where $J\big{(}\mathbf{s}_{k}\big{)}$ denotes the $(2N+1)^{2}\times 3$ Jacobian matrix of the warped intensity vector $\mathbf{t}_{F}\big{(}\mathbf{s}_{k}\big{)}$, evaluated at the nominal parameter values $\mathbf{s}_{k}$, and $\Delta\mathbf{s}_{k+1}$ denotes the perturbations of the parameters of the state vector. By substituting this approximation into Eq. (12) we get the following simpler objective function:
@@ 4.3 Lightweight event-ECC
To this end, consider the set $\mathcal{N}^{\prime}_{k+1}$ with the coordinates of the four template’s pixels whose densities were updated after the calculation of the warp, defined in Eq. (17). That is, this set contains the recently modified template’s pixels whose densities will be used in the next step of the tracking algorithm. Accordingly, we define the sets $\mathcal{N}^{\prime}_{{k+1}_{x}}$ $\mathcal{N}^{\prime}_{{k+1}_{y}}$ of the template’s gradient whose values have changed. Depending on the position of the pixels of $\mathcal{N}^{\prime}_{k+1}$ in the support area $\mathcal{S}_{T}$ of the template, the cardinality of the union $\mathcal{S}=\mathcal{N}^{\prime}_{{k+1}_{x}}\bigcup\mathcal{N}^{\prime}_{{k+1}_{y}}$ is bounded by $12$, that is $|\mathcal{S}|\leq~{}12$. Having defined the set $\mathcal{S}$ we can define its complement w.r.t. the set $\mathcal{S}_{\mathcal{T}_{F}}$, that is:
$\mathcal{S}^{c}=\mathcal{S}_{\mathcal{T}_{F}}\backslash\mathcal{S}$
@@
$J\big{(}\mathbf{s}_{k+1}\big{)}$ $=$ $\left[J_{\mathcal{S}^{c}}^{t}\big{(}\mathbf{s}_{k}\big{)}~{}J_{\mathcal{S}}^{t}\big{(}\mathbf{s}_{k+1}\big{)}\right]^{t}$ (18)
@@
$\mathbf{p}_{\mathbf{t}_{F_{\mathcal{S}}}}(\mathbf{s}_{k+1})$ $=$ $\mathbf{p}_{\mathbf{t}_{F_{\mathcal{S}}}}(\mathbf{s}_{k})+\Delta\mathbf{p}_{k+1}-\Delta\mathbf{p}_{k}$
$\mathbf{p}_{\mathbf{t}_{F}}(\mathbf{s}_{k+1})$ $=$ $\left[\mathbf{p}_{\mathbf{t}_{F_{\mathcal{S}^{c}}}}^{t}(\mathbf{s}_{k})~{}~{}\mathbf{p}_{\mathbf{t}_{F_{\mathcal{S}}}}^{t}(\mathbf{s}_{k+1})\right]^{t}$ (19)
@@
         Form the model’s vector $\mathbf{m}_{F}(T_{k+1})$ using Eqs. (5,6) and the updated set $\mathcal{E}_{\mathcal{M}}(T_{k+1})$
         Using Eq. (17), Compute the optimal warp $\Delta\mathbf{s}^{\star}_{k+1}$
@@
         Use $\boldsymbol{s}_{k+1}$ in Eq. (10) to find out where the central event $\mathbf{x}_{c}^{\prime}(T_{k+1})$ of set $\mathcal{E}_{\mathcal{M}}(k+1)$ is mapped
         Use the interpolation scheme (Eq. (6)) to Update the densities of template’s pixels $\in\mathcal{N}^{\prime}_{k+1}$
         Form the vector $\mathbf{t}_{F}(\mathbf{s}_{k+1})$ and the jacobian $J(\mathbf{s}_{k+1})$ according to Eq. (18)
         Update the quantities of Eq. (19)
         Update the tracking states $S_{F}=[S_{F}~{}\boldsymbol{s}_{k+1}]$